1.The height h of a cylinder equals the circumference of the cylinder.In terms of h,what is the volume of the cylinder?
(A)h^3/4 pi
(B)h^2/2 pi
(C)h^3/2
(D)h^2+4 pi
(E)pi h^3
Please help and thank you so much for giving your time to help me with this problem.

Answer by lwsshak3(11628)

1.The height h of a cylinder equals the circumference of the cylinder.In terms of h,what is the volume of the cylinder?
(A)h^3/4 pi
(B)h^2/2 pi
(C)h^3/2
(D)h^2+4 pi
(E)pi h^3
**
Volume of a cylinder=area of circular base*height
given circumference=h=2πr
r=h/2π
Area of circular base=πr^2=π(h/2π)^2=πh^2/4π^2=h^2/4π
Volume of cylinder=h*h^2/4π=h^3/4π
Ans: (A)